In a market with a duopoly, if market demand is P = 300 - Q find the Cournot Reaction curves and the Cournot quantity solutions then deduce the price in the case where Marginal cost curves for either of the duopoly firms is Mc = 91 +30 and Mc2 = 42 +30 . Compare your results to the case where a Monopolist that has a MC=Q +30 replaces the duopoly. What are the monopoly quantity and price? Which quantities are bigger, Cournot or Monopoly? What is the consumer Surplus in both cases? Set up the oligopoly model in a game-theoretical prisoner's dilemma framework. Explain briefly the strategies and how you reach the Nash Equilibrium.
Answer #1

Demand: P - 300.9,92 Marginal revenue : MR, = 300 -29,-902 MR,= 300 -9,- 292 Equilibrium: MR, = Ma . or, 300 - 29,-9, = 9,+ 3Oligopoly consumer swepuus: “So x ( 30o - 16 ) x 13 5 = 425 Monopol, Consumer sween: com : 5*( 300 - 210)*90 : 4050 In game fThen the residual demand for defect firm P = 300- 45 - qi - 255 - qs MR;= 255-298 mci = 30 + qs Equilibrium: MR j = MC j. ar